# Family Matters: Language Transfer and Merging for Adapting Small LLMs to Faroese

### v1

@@ 4.3 RQ2: Effects of Merging
We see in Table 3(b) that merging all transfer languages (setups Mergeeq and Mergeis+) leads to the lowest perplexities in full-parameter fine-tuning, suggesting that with sufficient learning capacity — and at least for language modelling — models can benefit from a mixture of related languages. For LoRA, however, the opposite holds: the fewer languages merged, the better the results, with the setups without merging performing best. Full fine-tuning appears more robust to language merging. Merging LoRA-fine-tuned models may generally be more problematic than merging fully fine-tuned models, as other work has also shown, possibly due to a weaker degree of representation alignment (Stoica et al., 2024). Zero-shot perplexities (Table 3(a)) in particular for Mergeeq are exploded; the more is merged the higher the perplexities. But this, interestingly, still results in a better initialization for full fine-tuning.
@@
In Table 1(f), we see that merges are beneficial for the full tuning setup: For the 135M model, the best 3 models are the 3 merges. For the 360M model, Icelandic is best, but followed by 2 merges. For LoRA however, the situation is different: For the 135M model, merges are within the same range as models with individual languages, while for the 360M models, the scores of merges are lower than for 3 out of 5 single languages.
@@
The Merge2 setup, which combines Icelandic and Danish equally, has four wins in Table 2(f) and is clearly the best of the merges. Full fine-tuning benefits particularly the Merge2 setup, which is best overall setup for 135M and runner-up for 360M. For LoRA, single-language setups however outperform merges. It is better than Icelandic only for both full fine-tuning models, while Icelandic only is better for both LoRA models, again demonstrating that merges show promise for full fine-tuning but are less suitable for LoRAs.
@@
In the preceding paragraphs, we found that merges are sometimes competitive but most often outperformed by individual languages: We found that Icelandic tends to give stronger gains on the linguistic probes in FoBLiMP, while Danish seems more helpful for comprehension probes in FoBCoMP. This raises the question of whether the two languages can complement each other, with Icelandic adding linguistic strengths and Danish adding text comprehension skills. We therefore ask: Do merges add linguistic capabilities to transfer from Danish only? To test this, we compare Danish-only with merges on FoBLiMP (Table 1). Comparing each merge pairwise against Danish within otherwise equal setups, the merge wins in 30/60 cases and Danish in 29 cases, and in one case, there is a tie. To see the potential of merging, we also compare the best merge against Danish. The best merge outperforms Danish in 14/20 cases, while Danish performs better in 6 cases. These results suggest that merging can indeed add linguistic capabilities to Danish models. Exploring this further is a promising direction for future work.
Finally, we ask: Do merges add comprehension capabilities to Icelandic models? For FoBCoMP, we make a pairwise comparison of Icelandic with the merged models (based on Table 2). Overall, the merges do not perform better: Icelandic wins in 38 out of 60 comparisons, while the merges win 16 times, with 6 ties. Focusing only on the best merge in each case, the results are evenly split, with 10 wins for the merge and 10 wins for Icelandic. This suggests that merges can sometimes match or improve on Icelandic only, but the potential is smaller than for linguistic capabilities.
@@
Model Overall Lexical Grammar Semantics Fluency LoRA-Mergeis+ $3.436$ $3.865$ $3.645$ $3.015$ $3.216$ LoRA-Is+ $3.396$ $3.983$ $3.587$ $2.891$ $3.126$ Full-Is+ $3.413$ $3.940$ $3.600$ $2.955$ $3.160$ Full-Mergeis+ $3.361$ $3.924$ $3.535$ $2.934$ $3.051$
Table 4: Sentence Continuation scores (averages over both annotators). Scale 0–5, higher is better. Pearson’s correlation coefficient between annotators: $r=0.546$.
Model Task Completion Linguistic Quality Full-Is+ $3.01$ $4.11$ Full-Mergeis+ $2.92$ $3.82$ LoRA-Is+ $1.20$ $2.02$ LoRA-Mergeis+ $1.03$ $1.78$
Table 5: Summarization scores (average over both annotators). Scale 0–5, higher is better. Pearson correlation coefficient between annotators: $r=0.879$.
@@ 4.4 RQ3: Full Fine-Tuning versus LoRA
Overall, full fine-tuning yields higher results than LoRA across most evaluations, but particularly in merging setups (as discussed in 4.3). This advantage is especially visible in the text comprehension tasks. LoRA shows its relative strength in linguistic acceptability transfer, especially for 135M models.
@@
In Table 1(f), we see that for the smaller 135M model, LoRA outperforms full fine-tuning in 5 out of 6 cases for single languages, while for the 360M model the results are balanced (3 wins each). Looking at all individual tasks in Figure 4, LoRA again has a clear advantage for the 135M model (34 wins vs. 23 for full fine-tuning, with 3 ties), but for the 360M model both methods perform similarly, with full tuning having a slight edge. For merges, however, both the mean scores in Table 1(f) and the win rates in Figure 4 indicate that full fine-tuning is better overall. These findings highlight that LoRA performs surprisingly well for the 135M model, capturing syntactic transfer efficiently despite being more parameter-efficient. Interestingly, this contradicts our initial assumption that higher learning capacity is more important for acquiring linguistic skills, while preventing catastrophic forgetting would primarily benefit comprehension tasks.
Figure 4: Wins of LoRA vs. full FT for model twins on FoBLiMP. No merges (left), merges (right).
@@
results for full tuning and LoRA are compared in Figure 5. Full tuning consistently outperforms LoRA: full tuning achieves 35 wins compared to 22 for LoRA (with 3 ties; see Figure 5). This may seem counterintuitive, as fine-grained language skills are not obviously critical for comprehension tasks. However, it is possible that full tuning provides a more stable adaptation process, which benefits performance on these datasets.
Figure 5: Wins of LoRA vs. full FT for model twins in FoBCoMP. No merges (left), merges (right).
Model size effects
Differences between full tuning and LoRA are mainly driven by the 135M model. For the 360M model, margins between both approaches are small, suggesting that larger models are more robust to adaptation method choice.
@@ 4.5 Qualitative Observations
The low semantic scores across models highlight a limitation in adaptation for low-resource languages. While the models acquire competency in surface-level linguistics — producing valid Faroese vocabulary, maintaining grammatical structures, and achieving natural-sounding fluency — they struggle significantly with generating meaningful, coherent content. This reflects a fundamental challenge in language model training: while syntactic and lexical patterns can be learned from limited data through transfer from related languages, semantic understanding — which requires world knowledge — is a challenge, particularly in small models. The example in Figure 6 highlights this: although the models produce text that appears fluent, they often fail to convey coherent ideas, limiting their utility for applications requiring content generation.
Many summarization outputs exhibit mixes between languages, as in the example in Figure 7. This was not the case for sentence continuation outputs, which indicates that the low-resource tuning destroyed some of the models’ linguistic abilities.
Faroese: Translation:
Input: Børn millum fimm og 11 ár verða nú koppsett fyri koronu nú Input: Children between five and 11 years are now being vaccinated for coronavirus now
Continuation: á døgninum. Síðani frísparksmeginreglurnar vórðu settar í gildi 24. apríl hevur eingin smitta verið staðfest, síðan stigtakarin til spæliskeiið - Danmark - varð settur at hava tamarhald á korona-smittuni týsdag Continuation: during the day. Since the free kick regulations were put into effect on April 24th, no infection has been confirmed, since the initiator of the play workshop - Denmark - was set to have control over the corona infection on Tuesday
Figure 6: Example: Semantic nonsense in outputs.
@@
4.6 Limitations and Future Work
The small amount of evaluation data for Faroese is a major limitation, and even the human study covers relatively few samples. Summarization results in particular should be viewed with caution as the small validation set made hyperparameter tuning difficult, and in particular the LoRA models may not have been trained optimally.
Valuable directions for future work are to investigate how these findings change with model sizes and after instruction tuning, and if adaptive or data-driven merging strategies can lead to better transfer from multiple languages. Multilingual studies could reveal whether the same patterns hold for other low-resource languages in other families.
@@ 5 Conclusion
We studied how to adapt small LLMs to Faroese by transferring from related Scandinavian languages, merging models trained on these languages, and comparing full and parameter-efficient fine-tuning. We found that transfer from related languages is essential, but the best source varies. Icelandic gave the strongest gains in linguistic accuracy, while Danish was more helpful for comprehension. This suggests that practical applications should draw on multiple sources rather than relying only on the closest relative. Merging showed potential, although the benefits were inconsistent and more sensitive under LoRA than full fine-tuning. This indicates that merging can add complementary strengths, but requires careful design. LoRA proved effective for improving linguistic acceptability, while full fine-tuning performed better on comprehension-heavy tasks. Full fine-tuning also provided a stronger base for downstream fine-tuning on summarization in low-resource settings. Overall, our results show that transfer from related languages is key for adapting LLMs to low-resource settings, and that the optimal tuning strategy depends on priorities and target application.

### v2

@@ 4.3 RQ2: Effects of Merging
We see in Table 11(b) that merging all transfer languages (setups Mergeeq and Mergeis+) leads to the lowest perplexities in full-parameter fine-tuning, suggesting that with sufficient learning capacity — and at least for language modelling — models can benefit from a mixture of related languages. For LoRA, however, the opposite holds: the fewer languages merged, the better the results, with the setups without merging performing best. Full fine-tuning appears more robust to language merging. Merging LoRA-fine-tuned models may generally be more problematic than merging fully fine-tuned models, as other work has also shown, possibly due to a weaker degree of representation alignment (Stoica et al., 2024). Zero-shot perplexities before adaptation on Faroese (Table 11(a); Appendix C) are extremely high for merges, in particular for Mergeeq; the more is merged, the higher the perplexities. But this, interestingly, still results in a better initialization for full fine-tuning.
@@
In Table 3(f), we see that merges are beneficial for the full tuning setup: For the 135M model, the best 3 models are the 3 merges. For the 360M model, Icelandic is best, but followed by 2 merges. For LoRA however, the situation is different: For the 135M model, merges are within the same range as models with individual languages, while for the 360M models, the scores of merges are lower than for 3 out of 5 individual languages.
@@
The Mergeda+is setup, which combines Icelandic and Danish equally, has four wins in Table 4(f) and is clearly the best of the merges. Full fine-tuning benefits particularly the Mergeda+is setup, which is best overall setup for 135M and runner-up for 360M. For LoRA, single-language setups however outperform merges. It is better than Icelandic only for both full fine-tuning models, while Icelandic only is better for both LoRA models, again demonstrating that merges show promise for full fine-tuning but are less suitable for LoRAs.
@@
Our preceding analyses indicate that merges are sometimes competitive but often outperformed by individual languages. Icelandic tends to drive stronger gains on linguistic probes (FoBLiMP), whereas Danish appears more beneficial for comprehension probes (FoBCoMP). This suggests that merges could potentially leverage the linguistic strengths of Icelandic while incorporating the comprehension skills of Danish. We find that merges can indeed add linguistic capabilities to Danish models. Comparing Danish-only models with merges on FoBLiMP (Table 3), pairwise comparisons show that merges outperform Danish in 30 out of 60 cases, Danish wins in 29 cases, and one case ties. Considering the best merge against Danish, the merge achieves 14 wins out of 20, compared to 6 for Danish. These results indicate that merging can meaningfully enhance the linguistic capabilities of Danish models. For comprehension capabilities, merges provide only limited benefits to Icelandic models. On FoBCoMP (Table 4), in 60 pairwise comparisons, Icelandic wins 38 times, merges 16, with 6 ties. Focusing only on the best merge per case, results are evenly split (10 wins each). This suggests that while merges can occasionally match or slightly improve comprehension over Icelandic-only models, the gains are smaller than those observed for linguistic capabilities.
@@
Overall Lex. Gram. Sem. Fluency LoRA-Mergeis+ 3.436 3.865 3.645 3.015 3.216 LoRA-Is 3.396 3.983 3.587 2.891 3.126 Full-Is 3.413 3.940 3.600 2.955 3.160 Full-Mergeis+ 3.361 3.924 3.535 2.934 3.051
(a) Sentence Continuation. Pearson $r=0.546$.
Task Compl. Ling. Quality 3.01 4.11 2.92 3.82 1.20 2.02 1.03 1.78
(b) Summarization. $r=0.879$.
Table 6: Results for the human evaluation. Scores are averages over annotators. Scale 0–5, higher is better.
Discussion
Our results confirm the observation from prior work that LoRAs are more challenging to merge than fully fine-tuned models (Stoica et al., 2024; Tang et al., 2024; Zhang and Zhou, 2025), so this outcome is consistent with expectations. More notably, we show that merging models fine-tuned on related languages can meaningfully benefit CPT in a new target language. Unlike task merging, where performance on the original tasks is generally preserved or improved (Yadav et al., 2024), previous attempts at language merging have shown that out-of-the-box performance often deteriorates (Glocker et al., 2025). [^5] Our out-of-the-box perplexities on Faroese FineWeb confirm this trend, with merged models performing worse than individual source languages; see Table 11(a) in Appendix C. In contrast, by continuing pre-training on the target language after merging, we observe gains especially in perplexity, demonstrating that while it may not lead to strong performance out-of-the-box, language merging is a very promising initialization strategy for further fine-tuning. We speculate that the merges lead to parameter interference, which the continued training helps resolve. Finally, although prior work indicates that larger models are generally easier to merge (Dang et al., 2024; Yadav et al., 2024), we do not observe a strong size effect in our CPT settings. This however does not preclude effects for substantially larger models, where merging may become even more impactful. Overall, our findings show that initializing with merges can leverage complementary capabilities across related languages, at least in the context of full fine-tuning.
@@ 4.4 RQ3: Full Fine-Tuning versus LoRA
In the mean results in Table 3(f), we observe that for the smaller 135M model, LoRA outperforms full fine-tuning in 5 out of 6 single-language settings. For the 360M model, however, performance is evenly split, with three wins each. Considering win rates across all individual tasks yields a similar pattern. For the 135M model, LoRA shows a clear advantage (22 wins vs. 7 for full fine-tuning, with 3 ties). In contrast, for the 360M model the two methods perform comparably, with a slight edge for full fine-tuning (12 wins for LoRA vs. 16 for full fine-tuning, with 2 ties). For merged transfer setups, both the mean scores in Table 3(f) and the overall win rates indicate an advantage for full fine-tuning. Across these conditions, LoRA achieves 8 wins, compared to 22 for full fine-tuning. These findings highlight that LoRA performs surprisingly well for individial languages in the 135M model. It captures syntactic transfer efficiently despite its parameter-efficiency. This contradicts our initial assumption that higher learning capacity is more important for acquiring linguistic skills, while preventing catastrophic forgetting would primarily benefit comprehension tasks.
@@
Overall, full fine-tuning consistently outperforms LoRA, achieving 35 wins compared to 22 for LoRA (with 3 ties). At first glance, this may seem somewhat unexpected, as fine-grained linguistic adaptation is not obviously more crucial for text comprehension tasks than for acceptability judgments. However, full fine-tuning may enable a more thorough and stable adaptation of the model, which in turn benefits performance on these datasets. The overall difference is primarily driven by the 135M model. Here, full fine-tuning clearly dominates (22 wins vs. 6 for LoRA, with 2 ties). For the 360M model, by contrast, the margin is small (16 wins for LoRA vs. 13 for full fine-tuning, with 1 tie), suggesting that the larger models are less sensitive to the choice of adaptation method and generally more robust across setups.
@@
Discussion
Our findings reveal a clear task-dependent pattern. Full fine-tuning is overall the more stable and robust adaptation method. Its advantage is most evident in perplexity reduction, text comprehension, and especially summarization. It achieves higher results than LoRA across most evaluations, but particularly in merging setups (as discussed in 4.3). LoRA shows its strength in linguistic acceptability, indicating that parameter-efficient adaptation is sufficient — and sometimes advantageous — for capturing syntactic regularities. This contrasts with our initial assumption that higher learning capacity would be most important for fine-grained and language-dependent linguistic skills: In prior work, LoRA is often associated with improved generalization and reduced catastrophic forgetting (Biderman et al., 2024; Ghosh et al., 2024). From this perspective, one might expect LoRA to be more beneficial for comprehension tasks, where preserving broader and more language-independent knowledge should matter most. Instead, we observe that full fine-tuning dominates in comprehension and summarization, while LoRA’s advantage is in linguistic acceptability.
Previous studies report that LoRA is particularly effective and stable in low-data regimes (Whitehouse et al., 2024; Vu et al., 2022). Although Faroese is a low-resource language, our CPT uses a 95M-word corpus, which does not clearly fall into the low-data scenarios typically examined in previous literature that looked at task adaptation, where low-data often means samples in the hundreds. Nevertheless, the gains for LoRA on acceptability tasks suggest that parameter-efficient adaptation helps to better learn core syntactic regularities and to transfer them to related languages.
@@ 4.5 Qualitative Observations
The low semantic scores across models confirm a limitation in small models, especially for low-resource languages like Faroese. While the models acquire competency in surface-level linguistics — producing valid Faroese vocabulary, maintaining grammatical structures, and achieving natural-sounding fluency — they struggle significantly with generating meaningful, coherent content. This reflects a fundamental challenge in language model training: while syntactic and lexical patterns can be learned from limited data through transfer from related languages, semantic understanding, which requires world knowledge, is a challenge.
Many summarization outputs exhibit mixes between languages, as in the example in Figure 2. This was not the case for sentence continuation outputs, which indicates that the low-resource tuning interferes with the models’ linguistic abilities.
@@
4.6 Generality of Results.
To assess generalizability, we replicate our approach for Northern Sámi (Uralic) and Upper Sorbian (West Slavic), both even lower-resourced than Faroese and less related to English, the language of the base model. The core findings hold: typologically more related languages (Finnish/Estonian for Northern Sámi; Czech/Polish for Upper Sorbian) consistently outperform majority/contact languages (Swedish/Norwegian; German) on perplexity, confirming our finding from Section 4.2 about typological similarity as an important factor for transfer effectiveness. As in Section 4.3, merged models under full fine-tuning again achieve the lowest perplexities, confirming merging as a particularly successful initialization strategy for further continued pre-training on the target language. On linguistic acceptability, LoRA scores are even here consistently better than full fine-tuning scores, confirming the result we found surprising, that LoRA is particularly good for linguistic acceptability, from Section 4.4. One notable divergence to the Faroese results is that LoRA outperforms full fine-tuning on perplexity for these two languages, possibly reflecting stronger regularization benefits when adapting to smaller or noisier corpora. Comprehension task results remain largely inconclusive across both languages. Details on the setup and full results are reported in Appendix D.
@@ 5 Conclusion
This article examined strategies for adapting small LLMs to Faroese, comparing transfer from related Scandinavian languages, model merging across multiple source languages, and full versus parameter-efficient LoRA fine-tuning. Our results establish that transfer from related languages is essential, but no single source language dominates: Icelandic proved most valuable for linguistic accuracy, while Danish contributed more to comprehension. This complementarity suggests that one should draw on multiple source languages simultaneously, which motivates experiments on model merging. Under full fine-tuning, merging consistently improved perplexity, showing that merged models are a successful initialization strategy for continued pre-training on Faroese. Gains were however less reliable for finer-grained probes, and merged LoRA adapters showed lower stability. The choice of adaptation strategy entails a trade-off: LoRA was more effective for linguistic acceptability judgments, whereas full fine-tuning produced stronger comprehension performance and a more robust foundation for downstream tasks.
Future work should examine if our results transfer to larger and instruction-tuned models, and explore whether adaptive or data-driven merging strategies can yield more consistent transfer gains.
Limitations
The scarcity of Faroese evaluation data is a fundamental limitation, both for this study and for low-resource NLP more broadly. Even our human evaluation covers relatively few samples, as linguist-administered rating is labor-intensive and costly. Summarization results in particular should be interpreted with caution: the small validation set was a much less reliable signal for hyperparameter tuning than a validation set of more significant size would be, and in particular the LoRA models, which are more sensitive to hyperparameters, may consequently not have been trained optimally.
We experiment with only two model sizes, one LoRA configuration, and restrict ourselves to English-only base models, chosen as the cleanest experimental baseline given the absence of small (<<1B) multilingual models with well-documented training data composition. Extending this work to multilingual base models that include (some of) the target languages would be a valuable next step.
The English-centric tokenizer of our base models is another limitation: it increases inference cost for non-English languages and may also affect model performance. For real-world deployment and to optimize performance, transferring to a multilingual tokenizer, or one specifically adapted to the Northern Germanic languages, would be preferable.
@@ Appendix A Training Details
This section provides additional details to support reproducibility, supplementing Section 3.1 and Section 3.4. Hyperparameters for continued pre-training are reported in Table 7; those for summarization fine-tuning are reported in Table 8.
@@
LoRA modules Feed-forward layers (following Kunz (2025))
@@
Appendix B Dataset Statistics
This section provides tabular overviews of our two new benchmark suites introduced in Section 3.3, detailing source datasets, subsets, modifications, and example counts. FoBLiMP is described in Table 9; FoBCoMP in Table 10. Both benchmarks are publicly released on HuggingFace under identifiers jekunz/FoBLiMP and jekunz/FoBCoMP, licensed under Creative Commons licenses: FoBLiMP under Creative Commons Attribution ShareAlike 4.0, FoBCoMP under Creative Commons Attribution Non-Commercial ShareAlike 4.0; consistent with the source datasets.
Corruption Type Source Subset Pairs Corruption Method
– Existing Work – – – Subject-Verb Agreement MultiBLiMP (Jumelet et al., 2025) Full 232 Automatic using morphological annotation
Verb Placement GermDetect (Michael and Horbach, 2025) All corrupted samples 2,026 Automatic using dependency trees
Flipped Words ScaLA (Nielsen, 2023) flip_neighbors 552 Automatic using simple heuristics
Deleted Words ScaLA (Nielsen, 2023) delete 601 Automatic using simple heuristics
Ours – Translation Pairs New Full 680 (from 100 source sentences) Machine-translated sentences ranked after human ratings
Table 9: Datasets and corruption methods used for the linguistic acceptability probes included in FoBLiMP.
Task Source Subset Pairs MinPair Generation
Sentiment Classification FoSent (Debess et al., 2024) Sentence; both annotators agree 91 Add a sentence at the end: This is good / bad.
Sentiment Classification FoSent (Debess et al., 2024) Document; both annotators agree 84 Add a sentence at the end: This is good / bad.
Topic Classification FoSent (Debess et al., 2024) Deemed relevant by human annotator 234 Add a sentence at the end: The topic of this article is: X. Exchange related topics (e.g., Local NewsInternational News).
Extractive QA FoQA (Simonsen et al., 2025) Full 21,867 (2,000 unique texts & correct answers) Dataset shuffling (negative answer passage is also part of the context, length controlled)
Extractive QA FoQA (Simonsen et al., 2025) Full 2,000 GPT-4-generated: extract a different, wrong answer with the same length.
Table 10: Datasets and corruption methods used for the text comprehension probes included in FoBCoMP.
Appendix C Extended Perplexity Results
Full LoRA 135M 360M 135M 360M En 73.61 58.27 73.61 58.27 +Da 50.98 40.45 63.23 44.38 +Is 38.77 30.09 40.05 30.54 +NoB 48.59 44.81 56.25 44.51 +NoN 61.58 50.28 68.02 56.62 +Sv 69.96 57.02 78.22 61.43 Mergeeq 103.45 42.40 182.21 338.30 Mergeis+ 94.95 40.16 68.81 62.88 Mergeda+is 65.28 40.39 78.46 146.20
(a) Before continuing training on Faroese.
Full LoRA 135M 360M 135M 360M 4.98 3.75 5.51 4.48 4.19 3.56 4.25 3.55 4.44 3.48 4.53 3.53 4.22 3.63 4.26 3.56 4.60 3.66 4.90 4.08 4.26 3.60 4.21 3.58 4.08 3.41 4.61 3.93 4.08 3.41 4.58 3.77 4.22 3.49 4.56 3.80
(b) After continuing training on Faroese.
Table 11: Average per-token perplexity on the Fineweb-2 evaluation set, extending results in Table 5.
See Table 11(a).
Appendix D Generality of the Results: Experiments on Northern Sámi and Upper Sorbian
To assess which of our Faroese findings generalize to other languages, we replicate the approach for two additional low-resource languages: Northern Sámi and Upper Sorbian. Both have fewer speakers and smaller corpora than Faroese and are classified by UNESCO as endangered (Northern Sámi) and vulnerable (Upper Sorbian) (Moseley, 2010). Neither functions as the majority language in its region, and public life is dominated by majority languages (Dagsvold et al., 2015; McMonagle, 2022; Mitschke, 2025). Typologically, Upper Sorbian is Indo-European (West Slavic), while Northern Sámi belongs to the Uralic family. This allows us to test whether the Faroese findings generalize both within Indo-European and across language families. [^6] Unlike the Faroese–Icelandic case, the closest transfer languages for Northern Sámi (Finnish) and Upper Sorbian (Czech and Polish) are not themselves low-resource, so we cannot test whether a more distant but higher-resource language improves semantic performance.
D.1 Experimental Setup
We focus on the 360M parameter models and compare full fine-tuning and LoRA using the same hyperparameters as in the Faroese experiments. For each target language and adaptation method, we create two merges: one combining two languages (a typologically close language and the regional majority language), and one including one or two additional related languages. For Northern Sámi, we use Finnish and Estonian as the related languages (Kahn and Valijärvi, 2017), and Swedish and Norwegian (Bokmål) as majority, high-contact languages. Most Northern Sámi speakers are bilingual in one of these languages (Kahn and Valijärvi, 2017), which influences local varieties (Jokinen et al., 2016) and results in lexical borrowing (Pietikäinen, 2008, 2010). For Upper Sorbian, we select Czech and Polish as West Slavic neighbors and German as the regional majority language; all Sorbs speak German, often as their dominant language (Mitschke, 2025).
Evaluation relies on massively multilingual benchmarks due to the lack of suitable curated resources. We report perplexity on the respective FineWeb-2 test sets and evaluate linguistic acceptability with MultiBLiMP (Jumelet et al., 2025) (2536 pairs for Northern Sámi, 186 for Upper Sorbian) and reading comprehension with MultiWikiQA (Smart, 2025) (4146 examples for Northern Sámi, 4906 for Upper Sorbian). For MultiWikiQA, we follow the non-LLM negative sampling procedure used in the Faroese experiments.
D.2 Results and Conclusions
Table 12.
Perplexity $\downarrow$ MultiBLiMP $\uparrow$ MultiWikiQA $\uparrow$ Transfer Full LoRA Full LoRA Full LoRA Se En 6.5180 6.7996 0.8052 0.8111 0.7786 0.7742 +Fi 6.0851 5.1539 0.8222 0.8722 0.7735 0.7726 +Ee 6.2498 5.4192 0.8261 0.8600 0.7646 0.7820 +NoB 6.2706 5.2190 0.7973 0.8324 0.7742 0.7735 +Sv 6.3907 5.4526 0.7993 0.8360 0.7670 0.7800 +Fi+Sv 6.0597 5.5669 0.8285 0.8391 0.7752 0.7656 +All 5.8965 5.8592 0.8206 0.8324 0.7742 0.7677 Hsb En 7.3967 8.5332 0.7366 0.7527 0.7790 0.7729 +Cs 6.4403 5.8660 0.8011 0.8280 0.7831 0.7927 +De 6.9620 6.5548 0.7258 0.7903 0.7650 0.7725 +Pl 6.5469 5.8850 0.7527 0.8441 0.7811 0.7915 +De+Pl 6.4882 6.5572 0.7527 0.7957 0.7644 0.7707 +All 6.3261 6.3270 0.7581 0.8280 0.7748 0.7729
Table 12: Results for Northern Sámi and Upper Sorbian across CPT and merging configurations.
Effect of transfer languages.
A first continued pre-training step on a related language consistently improves perplexity compared to models without a transfer language. In both languages, typologically related transfer languages yield the largest gains: Finnish and Estonian outperform Norwegian Bokmål and Swedish for Northern Sámi, and Czech and Polish outperform German for Upper Sorbian. This replicates our Faroese finding that typological similarity predicts transfer effectiveness. Majority and contact languages also help, but generally less so. For linguistic acceptability probes (MultiBLiMP), typologically closer languages consistently improve over the English-only baseline. For more distant contact languages, LoRA models show consistent improvements over English while fully fine-tuned models do not, suggesting that stronger regularization helps when transfer languages are less related. Results on comprehension probes are less clear. For Northern Sámi, MultiWikiQA differences are marginal across setups. Upper Sorbian shows somewhat clearer improvements from related-language transfer, though gains remain small overall. Overall, these results confirm our Faroese findings from Section 4.2: transfer languages help, typological similarity matters, and morphological similarity is particularly important for linguistic acceptability.
Effect of merging.
Under full fine-tuning, merged models again achieve the lowest perplexities, confirming our finding from Section 4.3 that combining multiple transfer sources improves language modeling. Merged LoRA adapters, however, underperform the best single-language variants. For linguistic acceptability, results are mixed. For Northern Sámi MultiBLiMP under full fine-tuning, the Finnish–Swedish merge scores highest, while for Upper Sorbian a single language (Czech) wins. Under LoRA, a single language always scores highest (Finnish and Polish, respectively). For MultiWikiQA, merges and individual languages perform similarly for Northern Sámi under full fine-tuning, while merged LoRA adapters score lowest and Polish is the consistent winner for Upper Sorbian. We conclude that merging reliably helps language modeling and can help linguistic acceptability, but does not help comprehension. This is consistent with the Faroese results in Section 4.3.
LoRA versus full fine-tuning.
Unlike in the Faroese experiments (Section 4.4), where full fine-tuning slightly outperformed LoRA on perplexity, LoRA often achieves lower perplexity for Northern Sámi and Upper Sorbian. The gap is even larger for linguistic acceptability, where LoRA consistently outperforms full fine-tuning, especially for Upper Sorbian. This may reflect stronger regularization when adapting to smaller or noisier corpora. QA results remain inconclusive, particularly for Northern Sámi where differences across setups are minimal.